OpenStudy (ohmybookness):

How do you convert y+8=2(x-1)^2 into standard form? Can someone show me step by step?

OpenStudy (ohmybookness):

@phi can you help me out here?

OpenStudy (ohmybookness):

I know the coordinates of the vertex is (1,8)

OpenStudy (phi):

can you multiply out (x-1)(x-1) ?

OpenStudy (ohmybookness):

x^2-2x+1

OpenStudy (phi):

now multiply each term by 2 (we are expanding 2(x-1)^2 )

OpenStudy (phi):

you should get \[ 2x^2 -4x+2\] so now you have \[ y+8= 2x^2 -4x+2\] the last step is add -8 to both sides

OpenStudy (phi):

btw, the vertex is at (1,-8)

OpenStudy (ohmybookness):

oops my bad

OpenStudy (ohmybookness):

@phi is it y=2(x-1)^2-8

OpenStudy (jdoe0001):

note: \(\bf y+8=2(x-1)^2 \cong =2(x-1)^2-8\)

OpenStudy (phi):

that is "vertex form" standard form is ax^2+bx + c which we almost have (see post up above) y+8= 2x^2 -4x+2

OpenStudy (jdoe0001):

\(\bf y+8=2(x-1)^2 \cong y=2(x-1)^2-8\) that is

OpenStudy (ohmybookness):

wait so what i wrote was not standard form? huh. I thought it was

OpenStudy (phi):

what was the original question?

OpenStudy (ohmybookness):

to convert y+8=2(x-1)^2 into standard form

OpenStudy (phi):

ok, then you are almost done. so far you have y+8= 2x^2 -4x+2

OpenStudy (ohmybookness):

yeah

OpenStudy (phi):

add -8 to both sides

OpenStudy (phi):

y+8-8= 2x^2 -4x+2-8 now simplify to get "standard form"